f the oF scale is thrice the oC scale, what are the values of oF and oC? in 2 decimal places



1
Expert's answer
2021-08-16T08:32:52-0400

It is known that "\degree{F}=1.8\times\degree{C}+32." Therefore, this relationship can be represented as an equation:

"\frac{\degree{F}}{\degree{C}}=\frac{1.8\times\degree{C}+32}{\degree{C}}=3"

Solving this equation for oC, we get 26.67.

Fahrenheit scale is thrice the Celsius scale, hence "\degree{F}=26.67\times3=80.01"


Answer: 26.67oC; 80.01oF